Pleae answer with proper calculation and table line tqvm


Gates Technologies LTD. Makes and sells one product, which has the following standard variables production costs per unit Direct Material cost (2KG at $10 per kg) Direct Labour cost (4 hours at $ 15 per hour) Variables production overhead cost ($3 per labour hour) 20 60 12 The budgeted selling price per unit is $150, and the production and sales budget for the coming this year are as follows: Production in Units 120,000 Sales in Units 100,000 Theres is no opening inventory at the beginning of 2016. Budgeted fixed production overhead cost are $15 per unit, and they are absorbed based on a normal production level of 100,000 units per annum. The b udgeted non- production costs for the coming year ar e as followings Variables non-production overhead costs S8 per unit sold Fixed non-production overhead costs per annum $2,000,000 Required: a) Prepare two income statement for the year of 2016, one using absorption and the other using marginal costing systems b) Prepare a reconciliation statement reconciling the profits under marginal and absorption costing c) Briefly explain why a difference arises between profit figures when using marginal and absorption costing principles
Answer #1
a) INCOME STATEMENT-ABSORPTION COSTING
Sales (100000*$150) 15000000
Cost of goods sold:
Beginning inventory 0
Cost of goods manufactured:
Direct materials (120000*$20) 2400000
Direct labor (120000*$60) 7200000
Variable overhead (120000*$12) 1440000
Fixed overhead assigned (120000*15) 1800000
Manufacturing costs for the year 12840000
Total manufacturing costs 12840000
Ending inventory = 12840000*20000/120000 = 2140000
Cost of goods sold (Unadjusted) 10700000
Fixed overhead overabsorbed (1800000-1500000) -300000
Adjusted cost of goods sold 10400000
Gross profit 4600000
Operating expenses:
Non-production overhead costs = (100000*8+2000000) = 2800000
Net operating income 1800000
INCOME STATEMENT-MARGINAL COSTING
Sales (100000*$150) 15000000
Variable Cost of goods sold:
Beginning inventory 0
Cost of goods manufactured:
Direct materials (120000*$20) 2400000
Direct labor (120000*$60) 7200000
Variable overhead (120000*$12) 1440000
Variable Manufacturing costs for the year 11040000
Total manufacturing costs 11040000
Ending inventory = 11040000*20000/120000 = 1840000
Variable cost of goods sold 9200000
Variable non-production overhead costs (100000*8) 800000
Total variable expenses 10000000
Contribution margin 5000000
Fixed costs:
Manufacturing (100000*$15) 1500000
Non production costs 2000000 3500000
Net operating income 1500000
b) NOI as per absorption costing income statement 1800000
Less: Fixed costs deferred in ending inventory = 20000*15 = 300000
NOI as per variable costing income statement 1500000
c) The difference arises due to the inclusion of fixed manufacturing costs in the valuation of inventory
under the absorption costing system; while under the variable costing system only variable costs
are included.
This will result in deferring of OH costs to the next period or in carry over of fixed costs from the
previous year under absorption costing system. As the value of inventory differs, the NOI also
differs.